OpenStudy (anonymous):

A material decays at a rate of 1.1% per year. If you start with 250 grams of the material, in how many years will there be 100 grams remaining? Round to one decimal place. Which equation correctly models the situation?

OpenStudy (anonymous):

A. 250=100(.989)^t B.250=100(98.9)^t C.100=250(.989)^t D.100=250(98.9)^t

OpenStudy (tkhunny):

Where's the "start with" button? Discard A and B immediately. Decay requires a base less less than 1. Discard that silly D. I think it's done.